### v2

@@ 3.2 Overall Architecture
Figure 2 illustrates the overall architecture of our proposed MLICv2 and MLICv2+. MLICv2 preserves the fundamental structure of MLIC++ while introducing several key enhancements to improve compression efficiency, particularly at high bit rates. First, we replace residual blocks following downsampling and upsampling operations with lightweight token mixing blocks, which enhance representational capacity while reducing computational complexity. Second, we upgrade the multi-reference entropy model to MEMv2 by incorporating enhanced inter-slice and intra-slice context modules with context reweighting and advanced positional embedding to better capture latent dependencies. Third, we introduce a selective compression module $f_{s}$ that intelligently predicts and skips zero elements in latent representations, reducing both bit rate and arithmetic coding time without compromising reconstruction quality. To demonstrate the potential for input-specific optimization, we propose MLICv2+, which iteratively refines latent features $\boldsymbol{y}$ and side information $\boldsymbol{z}$ based on the rate-distortion characteristics of individual inputs, enabling tailored compression strategies for different content types.
@@
3.3 Efficient Simple Token Mixing Transform
As illustrated in Figure 6, MLIC++ exhibits performance degradation at high bit rates due to limited capacity of transform modules to retain sufficient information in latent representations [76]. The representational capacity is fundamentally influenced by the number of channels. As demonstrated by Ballé et al. [6], when channel numbers are limited, the performance gap between low- and high-capacity models is negligible at low bit rates. However, this gap becomes pronounced as bit rates increase, with higher-capacity models achieving significantly better reconstruction quality.
Simply increasing channel numbers often results in prohibitive computational complexity. A more practical approach is to enhance the architectural design itself to improve transform capacity without substantially increasing computational cost. Recent work has adopted transformer layers [46] for transform coding [45, 40, 88, 48], demonstrating remarkable performance improvements. However, existing techniques [45] can introduce up to twice the complexity of original residual blocks in the MLIC series.
Inspired by the elegance of MetaFormer [81], which consists of token mixing and channel-wise interaction components, we investigate its effectiveness for LIC. However, MetaFormer’s use of pooling for token mixing, while efficient, is suboptimal for compression tasks due to information loss. We propose a simple token mixing block as illustrated in Figure 2. For token mixing, we employ a depth-wise residual block (DepthRB) [32] for nonlinear embedding, depth-wise convolution for spatial token mixing, and a gate block [32] for efficient channel-wise interactions. LayerNorm (LN) [5] is adopted for stable training. Given input feature $\boldsymbol{\omega}$, the process is formulated as:
$\boldsymbol{\omega}$ $=\boldsymbol{\omega}+\textrm{Conv}1\times 1(\textrm{DWConv}5\times 5(\textrm{DepthRB}(\textrm{LN}(\boldsymbol{\omega})))),$ (7)
$\boldsymbol{\omega}$ $=\boldsymbol{\omega}+\textrm{Gate}(\textrm{LN}(\boldsymbol{\omega})),$
where $\boldsymbol{\omega}\in\mathbb{R}^{c\times h\times w}$ represents the input feature with $c$, $h$, $w$ denoting channels, height, and width respectively. Stacking two token mixing blocks creates deeper transforms with enhanced expressiveness. Compared to previous residual blocks [16] consisting of two $3\times 3$ convolutions, our token mixing block achieves lower complexity while delivering significantly better performance, especially at high bit rates. The complexity analysis reveals that stacking two simple token-mixing blocks reduces multiply-accumulate operations (MACs) by 7.38% at $c=192$ and by 8.87% at $c=320$, compared with stacking a single residual block [16, 34, 33].
@@ 3.4 Enhanced Multi-Reference Entropy Modeling
The enhanced multi-reference entropy model (MEMv2) is presented in Figure 2. To address the limitations of MLIC++, we introduce several key innovations: hyperprior-based global correlation prediction, context reweighting to handle varying channel importance, 2D Rotary Positional Embedding (RoPE) and guided selective compression for improved context modeling and entropy coding.
3.4.1 Hyperprior-Guided Global Correlation Prediction
In MLIC++ [33], global similarity between anchor and non-anchor parts of the previous slice is exploited to predict current slice similarities. However, for the first slice $\hat{\boldsymbol{y}}^{0}$, no such reference exists. The only available conditional information is the hyperprior $\hat{\boldsymbol{H}}$, obtained by upsampling side information $\hat{\boldsymbol{z}}$ extracted from latent representation $\boldsymbol{y}$. As illustrated in Figure 3, the hyperprior preserves global similarity patterns similar to the original image. We leverage this property by estimating global similarity between $\hat{\boldsymbol{y}}_{ac}^{0}$ and $\hat{\boldsymbol{y}}_{na}^{0}$ using the similarity between anchor and non-anchor parts of the hyperprior, $\hat{\boldsymbol{H}}_{ac}$ and $\hat{\boldsymbol{H}}_{na}$. Following MLIC++, we employ linear attention [60] for global context modeling and introduce a Gate block [32] for enhanced pointwise interaction:
@@ 3.4.1 Hyperprior-Guided Global Correlation Prediction
where $\hat{\boldsymbol{H}}_{na,q}=\text{Linear}(\hat{\boldsymbol{H}}_{na})$, $\hat{\boldsymbol{H}}_{ac,k}=\text{Linear}(\hat{\boldsymbol{H}}_{ac})$, and $\hat{\boldsymbol{y}}^{0}_{g,v}=\text{Linear}(\hat{\boldsymbol{y}}_{ac}^{0})$.
While prior methods [30, 34, 33, 32] capture only local dependencies in the first slice, our hyperprior-guided mechanism models global dependencies effectively. Given the information compaction property of latent representations [30], the fact that the first slice typically exhibits higher entropy as it serves as reference for subsequent slices, incorporating global context significantly reduces bit rate and improves compression performance.
@@ 3.4.2 Context Reweighting
In MLIC++ [33], the contexts are extracted by spatial attention mechanism. Given the query $\boldsymbol{Q}\in\mathbb{R}^{L\times c}$, key $\boldsymbol{K}\in\mathbb{R}^{L\times c}$ and the value $\boldsymbol{V}\in\mathbb{R}^{L\times c}$, where $L=hw$ is the sequence length, $h$ is the height, $w$ is the width, and $c$ is the channel number, in most spatial attention [70, 60], the attention map is computed by $\textrm{Sim}(\boldsymbol{Q}\boldsymbol{K}^{\top})\in\mathbb{R}^{L\times L}$, where Sim is the predefined function, such as Softmax [70] or two independent Softmax operations [60], indicating the information is gathered and weighted from spatial tokens ($\in\mathbb{R}^{1\times c}$). However, one common limitation is that all channel-wise features are treated equally, without considering that different channels may carry varying levels of importance. We argue that adaptively reweighting channel-wise context features can lead to more accurate entropy modeling. By suppressing noisy or less informative channels and emphasizing more relevant ones, the model has the potential to capture context information more precisely and improve overall compression performance, which has not been considered in existing methods [34, 33, 30, 31, 51, 19, 22, 54, 55].
@@
To address this issue, we introduce a two-stage context modeling strategy, where spatial aggregation [33] is first performed, followed by adaptive channel-wise context reweighting to capture the varying importance of different features. The architecture is in Fig. 4. Specifically, in context reweighting, the context is transformed to query $\boldsymbol{Q}$, key $\boldsymbol{K}$, and value $\boldsymbol{V}$, and the query and values are transposed to compute the channel-wise attention map [3]. The process is formulated as:
$\boldsymbol{O}=\textrm{Softmax}(\boldsymbol{Q}^{\top}\boldsymbol{K})\boldsymbol{V}^{\top}.$
@@
The resolution of the attention map $\textrm{Softmax}(\boldsymbol{Q}^{\top}\boldsymbol{K})$ is $c\times c$, indicating the importance of each channel ($\in\mathbb{R}^{L\times 1}$). The complexity of Eq. (8) is $\mathcal{O}(c^{2}hw)$. $\boldsymbol{O}$ is then fed into a Gate block [32] for efficient point-wise interactions. This attention-based reweighting helps the model focus on the most relevant parts of the context. To the best of our knowledge, this is the first attempt to explicitly model channel-wise importance in context features for LIC. Our approach brings consistent performance improvements with minimal additional cost.
3.4.3 2D Rotary Position Embedding
@@
MLIC++ [33] applies relative position embedding (RPE) from Swin-Transformer [46] directly to attention maps as additional bias. However, this direct addition may limit interaction with attention weights. Recently, Rotary Positional Embedding (RoPE) [63], a multiplication-based approach, has been widely adopted in language models [66, 58]. RoPE offers several advantages: compatibility with linear attention (whereas RPE requires adding biases that conflict with implicit attention computation), better extrapolation capabilities, and more concise implementation. For 2D image content, we extend the original 1D RoPE to handle horizontal, vertical, and diagonal relationships. Given query $\boldsymbol{q}_{m_{x},m_{y}}$ and key $\boldsymbol{k}_{n_{x},n_{y}}$ at 2D positions $(m_{x},m_{y})$ and $(n_{x},n_{y})$, we rotate queries and keys twice with different angles $\theta_{x}$ and $\theta_{y}$ for horizontal and vertical directions. When $c=2$, the rotation matrix for the query becomes:
@@
where $\theta_{(m_{x}-n_{x},m_{y}-n_{y})}=(m_{x}-n_{x})\theta_{x}+(m_{y}-n_{y})\theta_{y}$ encodes relative positions. Due to the ambiguity in defining relative importance of horizontal versus vertical positional dependencies, we make angles $\theta_{x}$ and $\theta_{y}$ learnable, initializing them to 10000 following original RoPE [63].
3.4.4 Guided Selective Compression
@@
Recent selective compression methods [39, 2] have employed neural networks to predict selective compression maps, but they are primarily applied to simpler entropy models. In addition, scale-based prediction has been introduced in recent works [62, 84]. However, when extended to advanced entropy models [34, 33], the training phase becomes highly unstable, with the overall loss value becoming “nan” after several steps, regardless of whether the parameters of the compression model are frozen or not. The observed training instability is primarily caused by the dependency introduced by the context modeling, rather than by the selective compression mechanism itself. Specifically, skipping non-zero elements in the current slice can significantly affect the entropy modeling of subsequent slices due to the inter-latent dependencies established by the context model. Considering the substantial performance gains enabled by context modeling, our goal is to skip the encoding of zero elements as much as possible—on top of the existing context modeling framework [51, 30, 33, 34]—in order to reduce the time spent on arithmetic encoding and decoding.
To address above issues, the guided selective compression module $f_{s}$ is proposed for post-training. This module is employed on the converged MLICv2, with the weights of MLICv2 frozen for stable training. The architecture of the guided selective compression $f_{s}$ module is presented in Fig. 2. Specifically, in $f_{s}$, the scale-based selective prediction is employed as the initial predicted selective compression map. For $j$-th element ${s}_{\boldsymbol{\sigma},j}$ in $\boldsymbol{s}_{\boldsymbol{\sigma}}$, the scale-based selective compression value is:
@@
where $\xi$ is a predefined threshold, set to $0.3$. For $\boldsymbol{y}^{i}_{ac}$, the decoded elements of previous slices serve as priors. For $\boldsymbol{y}^{i}_{na}$, the decoded elements of previous slices and $\boldsymbol{y}^{i}_{ac}$ serve as priors. The hyperprior $\hat{\boldsymbol{H}}$ also serves as a prior. The predicted selective map is formulated as:
$\boldsymbol{\varepsilon}^{i}_{ac}$ $=f_{s}\left(\boldsymbol{s}_{\boldsymbol{\sigma},ac}^{i},\lceil\boldsymbol{\Lambda}^{\textless i}\rfloor\odot\boldsymbol{s}^{\textless i},\hat{\boldsymbol{H}}\right),\boldsymbol{s}^{i}_{ac}=\textrm{STE}\left(\textrm{Sigmoid}\left(\boldsymbol{\varepsilon}^{i}_{ac}\right)\right),$ (12)
$\boldsymbol{\varepsilon}^{i}_{na}$ $=f_{s}\Bigl{(}\boldsymbol{s}_{\boldsymbol{\sigma},na}^{i},\lceil\boldsymbol{\Lambda}^{\textless i}\rfloor\odot\boldsymbol{s}^{\textless i},\lceil\boldsymbol{\Lambda}^{i}_{ac}\rfloor\odot\boldsymbol{s}^{i}_{ac},\hat{\boldsymbol{H}}\Bigr{)},\boldsymbol{s}^{i}_{na}=\textrm{STE}\left(\textrm{Sigmoid}\left(\boldsymbol{\varepsilon}^{i}_{na}\right)\right).$
During training, $\lceil\boldsymbol{\Lambda}\rfloor$ is known, therefore cross-entropy loss is employed to optimize $f_{s}$, treating selective prediction as a binary classification task. Additionally, elements with larger values have a higher impact on distortion, so the cross-entropy loss is weighted by $\lceil\boldsymbol{\Lambda}\rfloor+1$. To the best of our knowledge, we are the first to highlight the importance of quantization residuals and incorporate a weighted loss accordingly. The overall loss is:
@@
$\mathcal{L}_{s}=\sum_{i=0}^{L}\Bigl{(}(\lceil\boldsymbol{\Lambda}^{i}_{ac}\rfloor+1)\odot\mathcal{L}^{i}_{ce,ac}+(\lceil\boldsymbol{\Lambda}^{i}_{na}\rfloor+1)\odot\mathcal{L}^{i}_{ce,na}\Bigr{)},$ (13)
$\mathcal{L}_{ce,ac}^{i}=-\frac{\lceil\boldsymbol{\Lambda}^{i}_{ac}\rfloor\odot\log\boldsymbol{\varepsilon}^{i}_{ac}+\left(1-\lceil\boldsymbol{\Lambda}^{i}_{ac}\rfloor\right)\odot\log(1-\boldsymbol{\varepsilon}^{i}_{ac})}{2},$
$\mathcal{L}_{ce,na}^{i}=-\frac{\lceil\boldsymbol{\Lambda}^{i}_{na}\rfloor\odot\log\boldsymbol{\varepsilon}^{i}_{na}+\left(1-\lceil\boldsymbol{\Lambda}^{i}_{na}\rfloor\right)\odot\log(1-\boldsymbol{\varepsilon}^{i}_{na})}{2}.$
3.5 Iterative Latent Refinement
To validate our MLICv2’s potential and demonstrate achievable performance improvements, we employs encoding-time scaling to find optimal quantized latent representation $\hat{\boldsymbol{y}}$ and side information $\hat{\boldsymbol{z}}$ for each input. Direct encoder optimization requires storing model weights, gradients and intermediate features, leading to prohibitive memory consumption for high-resolution images. We adopt Stochastic Gumbel Annealing (SGA) [80] to refine $\boldsymbol{y}$ and $\boldsymbol{z}$ efficiently. In SGA, outputs of $g_{a}$ and $h_{a}$ serve as initialization. For anchor part $\boldsymbol{y}^{i}_{ac}$, $\boldsymbol{\Lambda}^{i}_{ac}=\boldsymbol{y}^{i}_{ac}-\boldsymbol{\mu}^{i}_{ac}$ is stochastically rounded up or down. The stochastic rounding direction is sampled from a relaxed one-hot categorical distribution:

### v3

@@ 3.2 Overall Architecture
Figure 3 illustrates the overall architecture of our proposed MLICv2 and MLICv2+. MLICv2 retains the fundamental design of MLIC++ but introduces three key enhancements to improve compression efficiency, especially at high bit rates. First, residual blocks of transforms are replaced with lightweight token mixing blocks, enhancing representational capacity while reducing computational cost. Second, the multi-reference entropy model is upgraded to MEMv2, which integrates improved inter-slice and intra-slice context modules with context reweighting and advanced positional embeddings to better capture latent dependencies. Third, a selective compression module $f_{s}$ is introduced to predict and skip zero elements in latent representations, effectively lowering both bit rate and arithmetic coding time. Furthermore, MLICv2+ extends MLICv2 by incorporating iterative refinement of latent features $\boldsymbol{y}$ and side information $\boldsymbol{z}$ according to the rate-distortion characteristics of each input, enabling adaptive and content-specific compression strategies.
3.3 Efficient Simple Token Mixing Transform (STMT)
As illustrated in Figure 7, MLIC++ suffers from performance degradation at high bit rates due to the limited capacity of its transform modules to retain sufficient information in latent representations xu2022multi. The representational capacity is largely determined by the number of channels. Ballé et al. balle2020nonlinear showed that when channel numbers are small, the performance gap between low- and high-capacity models is negligible at low bit rates but becomes significant as the bit rate increases, with higher-capacity models achieving noticeably better reconstruction quality.
Simply enlarging the channel dimension, however, leads to prohibitive computational costs. A more effective strategy is to enhance the transform design itself to improve capacity without substantially increasing complexity. Recent studies have explored transformer-based transforms liu2021swin, liu2023learned, li2024frequencyaware, zou2022the, lu2021tic, which achieve strong performance gains but often double the computational load compared to conventional residual blocks cheng2020learned, jiang2022mlic.
Inspired by the MetaFormer framework yu2022metaformer, which separates token mixing and channel interaction, we design a lightweight token mixing block tailored for learned image compression. Unlike MetaFormer’s pooling-based token mixing that causes information loss, our design adopts a depth-wise residual block (DepthRB) jiang2024llic for nonlinear embedding, a depth-wise convolution for spatial mixing, and a gate block jiang2024llic for efficient channel interactions. Layer normalization (LN) ba2016layer is applied for training stability. Given an input feature $\boldsymbol{\omega}\in\mathbb{R}^{c\times h\times w}$, the process is formulated as:
$\boldsymbol{\omega}$ $=\boldsymbol{\omega}+\mathrm{Conv}_{1\times 1}(\mathrm{DWConv}_{5\times 5}(\mathrm{DepthRB}(\mathrm{LN}(\boldsymbol{\omega})))),$ (6)
$\boldsymbol{\omega}$ $=\boldsymbol{\omega}+\mathrm{Gate}(\mathrm{LN}(\boldsymbol{\omega})).$
Stacking two token mixing blocks yields deeper transforms with enhanced expressiveness. Compared to conventional residual blocks with two $3\times 3$ convolutions cheng2020learned, our design achieves lower complexity and superior rate-distortion performance, particularly at high bit rates. Complexity analysis shows that stacking two token mixing blocks reduces MACs by 7.38% at $c=192$ and by 8.87% at $c=320$, compared to a single residual block cheng2020learned, jiang2023mlicpp.
@@ 3.4 Enhanced Multi-Reference Entropy Modeling
The enhanced multi-reference entropy model (MEMv2) is illustrated in Figure 3. To overcome the limitations of MLIC++, MEMv2 incorporates several key innovations: (i) hyperprior-guided global correlation prediction (HGCP) for more accurate dependency modeling, (ii) context reweighting (CR) to adaptively emphasize informative channels, (iii) 2D Rotary Positional Embedding (2D RoPE) for enhanced spatial awareness, and (iv) guided selective compression (GSC) to improve entropy coding efficiency. Together, these components enable more expressive and adaptive context modeling, particularly at high bit rates.
3.4.1 Hyperprior-Guided Global Correlation Prediction (HGCP)
In MLIC++ jiang2023mlicpp, global similarity between anchor and non-anchor parts of the previous slice is exploited to predict current slice similarities. However, for the first slice $\hat{\boldsymbol{y}}^{0}$, no such reference exists. The only available conditional information is the hyperprior $\hat{\boldsymbol{H}}$, obtained by upsampling side information $\hat{\boldsymbol{z}}$ extracted from latent representation $\boldsymbol{y}$. As illustrated in Figure 4, the hyperprior preserves global similarity patterns similar to the original image. We leverage this property by estimating global similarity between $\hat{\boldsymbol{y}}_{ac}^{0}$ and $\hat{\boldsymbol{y}}_{na}^{0}$ using the similarity between anchor and non-anchor parts of the hyperprior, $\hat{\boldsymbol{H}}_{ac}$ and $\hat{\boldsymbol{H}}_{na}$. Following MLIC++, we employ linear attention shen2021efficient for global context modeling and employ a Gate block introduced in LLIC jiang2024llic for enhanced pointwise interaction:
@@ 3.4.1 Hyperprior-Guided Global Correlation Prediction (HGCP)
where $\hat{\boldsymbol{H}}_{na,q}={\text{LinearLayer}}(\hat{\boldsymbol{H}}_{na})$, $\hat{\boldsymbol{H}}_{ac,k}={\text{LinearLayer}}(\hat{\boldsymbol{H}}_{ac})$, and $\hat{\boldsymbol{y}}^{0}_{g,v}={\text{LinearLayer}}(\hat{\boldsymbol{y}}_{ac}^{0})$.
While prior methods he2022elic, jiang2022mlic, jiang2023mlicpp, jiang2024llic capture only local dependencies in the first slice, our hyperprior-guided mechanism effectively models global dependencies. Due to the information compaction property of latent representations he2022elic, the first slice typically exhibits higher entropy as it serves as a reference for subsequent slices. Incorporating global context at this stage thus significantly reduces bit rate and enhances compression performance.
@@ 3.4.2 Context Reweighting (CR)
In MLIC++ jiang2023mlicpp, the contexts are extracted by spatial attention mechanism. Given the query $\boldsymbol{Q}\in\mathbb{R}^{L\times c}$, key $\boldsymbol{K}\in\mathbb{R}^{L\times c}$ and the value $\boldsymbol{V}\in\mathbb{R}^{L\times c}$, where $L=hw$ is the sequence length, $h$ is the height, $w$ is the width, and $c$ is the channel number, in most spatial attention vas2017attention, shen2021efficient, the attention map is computed by $\textrm{Sim}(\boldsymbol{Q}\boldsymbol{K}^{\top})\in\mathbb{R}^{L\times L}$, where Sim is the predefined function, such as Softmax vas2017attention or two independent Softmax operations shen2021efficient, indicating the information is gathered and weighted from spatial tokens ($\in\mathbb{R}^{1\times c}$). However, one common limitation is that all channel-wise features are treated equally, without considering that different channels may carry varying levels of importance. We argue that adaptively reweighting channel-wise context features can lead to more accurate entropy modeling. By suppressing noisy or less informative channels and emphasizing more relevant ones, the model has the potential to capture context information more precisely and improve overall compression performance, which has not been considered in existing methods jiang2022mlic, jiang2023mlicpp, he2022elic, he2021checkerboard, minnen2020channel, duan2023qarv, fu2023learned, qian2022entroformer, qian2020learning.
@@
To address this issue, we introduce a two-stage context modeling strategy, where spatial aggregation jiang2023mlicpp is first performed, followed by adaptive channel-wise context reweighting to capture the varying importance of different features. The architecture is in Fig. 5. Specifically, in context reweighting, the context is transformed to query $\boldsymbol{Q}$, key $\boldsymbol{K}$, and value $\boldsymbol{V}$, and the query and values are transposed to compute the channel-wise attention map ali2021xcit. The process is formulated as:
$\boldsymbol{O}=\textrm{Softmax}(\boldsymbol{Q}^{\top}\boldsymbol{K})\boldsymbol{V}^{\top}.$ (8)
@@
The attention map $\boldsymbol{M}=\textrm{Softmax}(\boldsymbol{Q}^{\top}\boldsymbol{K})$ is row-wise normalized via the Softmax function: $\boldsymbol{M}_{i,j}=\frac{\exp(\boldsymbol{q}_{i}^{\top}\boldsymbol{k}_{j})}{\sum_{j^{\ast}=1}^{C}\exp(\boldsymbol{q}_{i}^{\top}\boldsymbol{k}_{j^{\ast}})},$ where $\boldsymbol{q}_{i}$ and $\boldsymbol{k}_{j}$ are the $i$-th and $j$-th channel vectors of $\boldsymbol{Q}$ and $\boldsymbol{K}$, respectively. The summation is over spatial indices while channel indices remain open. The resolution of the attention map $\textrm{Softmax}(\boldsymbol{Q}^{\top}\boldsymbol{K})$ is $c\times c$, indicating the importance of each channel ($\in\mathbb{R}^{L\times 1}$). The complexity of Eq. (8) is $\mathcal{O}(c^{2}hw)$. $\boldsymbol{O}$ is then fed into a Gate block jiang2024llic for efficient point-wise interactions. This attention-based reweighting helps the model focus on the most relevant parts of the context. To the best of our knowledge, this is the first attempt to explicitly model channel-wise importance in context features for LIC. Our approach brings consistent performance improvements with minimal additional cost.
3.4.3 2D Rotary Position Embedding (2D RoPE)
@@
MLIC++ jiang2023mlicpp applies relative position embedding (RPE) from Swin-Transformer liu2021swin directly to attention maps as additional bias. However, this direct addition may limit interaction with attention weights. Recently, Rotary Positional Embedding (RoPE) su2024roformer, a multiplication-based approach, has been widely adopted in language models touvron2023llama, reid2024gemini. RoPE offers several advantages: compatibility with linear attention (whereas RPE requires adding biases that conflict with implicit attention computation), better extrapolation capabilities, and more concise implementation. For 2D image content, we extend the original 1D RoPE to handle horizontal (width), vertical (height), and diagonal relationships. Given query $\boldsymbol{q}_{m_{x},m_{y}}$ and key $\boldsymbol{k}_{n_{x},n_{y}}$ at 2D positions $(m_{x},m_{y})$ and $(n_{x},n_{y})$, we rotate queries and keys twice with different angles $\theta_{x}$ and $\theta_{y}$ for horizontal and vertical directions. When $c=2$, the rotation matrix for the query becomes:
@@
where $\theta_{(m_{x}-n_{x},m_{y}-n_{y})}=(m_{x}-n_{x})\theta_{x}+(m_{y}-n_{y})\theta_{y}$ encodes relative positions. Due to the ambiguity in defining relative importance of horizontal versus vertical positional dependencies, we make angles $\theta_{x}$ and $\theta_{y}$ learnable, initializing them to 10000 following original RoPE su2024roformer. For $c\textgreater 2$, the queries and keys are divided into $\frac{c}{2}$ groups, each following Equation (10).
3.4.4 Guided Selective Compression (GSC)
@@
Recent selective compression methods lee2022selective, alexandre2023hierarchical have employed neural networks to predict selective compression maps, but they are primarily applied to simpler entropy models. In addition, scale-based prediction has been introduced in recent works shi2022alphavc, zhang2024practical. However, when extended to advanced entropy models jiang2022mlic, jiang2023mlicpp, the training phase becomes highly unstable, with the overall loss value becoming “nan” after several steps, regardless of whether the parameters of the compression model are frozen or not. The observed training instability is primarily caused by the dependency introduced by the context modeling, rather than by the selective compression mechanism itself. Specifically, skipping non-zero elements in the current slice can significantly affect the entropy modeling of subsequent slices due to the inter-latent dependencies established by the context model. Considering the substantial performance gains enabled by context modeling, our goal is to skip the encoding of zero elements as much as possible—on top of the existing context modeling framework minnen2020channel, he2022elic, jiang2023mlicpp, jiang2022mlic—in order to reduce the time spent on arithmetic encoding and decoding.
To address above issues, the guided selective compression module $f_{s}$ is proposed for post-training. This module is employed on the converged MLICv2, with the weights of MLICv2 frozen for stable training. The architecture of the guided selective compression $f_{s}$ module is presented in Fig. 3. Specifically, in $f_{s}$, the scale-based selective prediction is employed as the initial predicted selective compression map. For $j$-th element ${s}_{\boldsymbol{\sigma},j}$ in $\boldsymbol{s}_{\boldsymbol{\sigma}}$, the scale-based selective compression value is served as the initialization:
@@
where $\xi$ is a predefined threshold, set to $0.3$. For $\boldsymbol{y}^{i}_{ac}$, the decoded elements of previous slices serve as priors. For $\boldsymbol{y}^{i}_{na}$, the decoded elements of previous slices and $\boldsymbol{y}^{i}_{ac}$ serve as priors. The hyperprior $\hat{\boldsymbol{H}}$ also serves as a prior. The scale-based initialization is refined by $f_{s}$ for final skip map:
$\boldsymbol{\varepsilon}^{i}_{ac}$ $={\textrm{Sigmoid}}\left(f_{s}\left(\boldsymbol{s}_{\boldsymbol{\sigma},ac}^{i},\lceil\boldsymbol{\Lambda}^{\textless i}\rfloor\odot\boldsymbol{s}^{\textless i},\hat{\boldsymbol{H}}\right)\right),\boldsymbol{s}^{i}_{ac}=\textrm{STE}\left(\boldsymbol{\varepsilon}^{i}_{ac}\right),$ (12)
$\boldsymbol{\varepsilon}^{i}_{na}$ $={\textrm{Sigmoid}}\left(f_{s}\left(\boldsymbol{s}_{\boldsymbol{\sigma},na}^{i},\lceil\boldsymbol{\Lambda}^{\textless i}\rfloor\odot\boldsymbol{s}^{\textless i},\lceil\boldsymbol{\Lambda}^{i}_{ac}\rfloor\odot\boldsymbol{s}^{i}_{ac},\hat{\boldsymbol{H}}\right)\right),\boldsymbol{s}^{i}_{na}=\textrm{STE}\left(\boldsymbol{\varepsilon}^{i}_{na}\right).$
During training, $\lceil\boldsymbol{\Lambda}\rfloor$ is known, therefore cross-entropy loss is employed to optimize $f_{s}$, treating selective prediction as a binary classification task. Additionally, elements with larger values have a higher impact on distortion, so the cross-entropy loss is weighted by $|\lceil\boldsymbol{\Lambda}\rfloor|+1$. The overall loss is:
@@
${\mathcal{L}_{s}=\sum_{i=0}^{L}\Bigl((|\lceil\boldsymbol{\Lambda}^{i}_{ac}\rfloor|+1)\odot\mathcal{L}^{i}_{ce,ac}+(|\lceil\boldsymbol{\Lambda}^{i}_{na}\rfloor|+1)\odot\mathcal{L}^{i}_{ce,na}\Bigr)},$ (13)
${\mathcal{L}_{ce,ac}^{i}=\tfrac{1}{2}\!\left[-\mathbf{1}\!\big(|\lceil\boldsymbol{\Lambda}^{i}_{ac}\rfloor|>0\big)\odot\log\boldsymbol{\varepsilon}^{i}_{ac}-\big(1-\mathbf{1}\!\big(|\lceil\boldsymbol{\Lambda}^{i}_{ac}\rfloor|>0\big)\big)\odot\log(1-\boldsymbol{\varepsilon}^{i}_{ac})\right]},$
${\mathcal{L}_{ce,na}^{i}=\tfrac{1}{2}\!\left[-\mathbf{1}\!\big(|\lceil\boldsymbol{\Lambda}^{i}_{na}\rfloor|>0\big)\odot\log\boldsymbol{\varepsilon}^{i}_{na}-\big(1-\mathbf{1}\!\big(|\lceil\boldsymbol{\Lambda}^{i}_{na}\rfloor|>0\big)\big)\odot\log(1-\boldsymbol{\varepsilon}^{i}_{na})\right]}.$
Since both the encoder and decoder compute identical selective compression maps using $f_{s}$ with the same inputs, the selective coding decisions are perfectly synchronized without requiring any additional information transmission or bitstream overhead. To the best of our knowledge, we are the first to highlight the importance of quantization residuals and incorporate a weighted loss accordingly.
3.5 Iterative Latent Refinement (ILR)
To validate our MLICv2’s potential and demonstrate achievable performance improvements, we employs encoding-time scaling to find optimal quantized latent representation $\hat{\boldsymbol{y}}$ and side information $\hat{\boldsymbol{z}}$ for each input. Direct encoder optimization requires storing model weights, gradients and intermediate features, leading to prohibitive memory consumption for high-resolution images. We adopt Stochastic Gumbel Annealing (SGA) yang2020improving to refine $\boldsymbol{y}$ and $\boldsymbol{z}$ efficiently. In SGA, outputs of $g_{a}$ and $h_{a}$ serve as initialization. For anchor part $\boldsymbol{y}^{i}_{ac}$, $\boldsymbol{\Lambda}^{i}_{ac}=\boldsymbol{y}^{i}_{ac}-\boldsymbol{\mu}^{i}_{ac}$ is stochastically rounded up or down. The stochastic rounding direction is sampled from a relaxed one-hot categorical distribution: